THE S_{1}(n, π^{∗}) STATE OF 2-CYCLOHEXEN-1-ONE: CAVITY RINGDOWN ABSORPTION SPECTRUM AND DFT CALCULATIONS

E. J. GILLES, L. K. AUSMAN, E. A. BROWN, \underline{S. DRUCKER}, Department of Chemistry, University of Wisconsin-Eau Claire, Eau Claire, WI 54702; S. LEE, J. CHOO, Department of Chemistry, Hanyang University, Ansan 425-791, Korea; M. RISHARD, J. LAANE, Department of Chemistry, Texas A & M University, College Station, TX 77843.

The cavity ringdown absorption spectra of 2-cyclohexen-1-one (2CHO) and a deuterated derivative were recorded near 380 nm in a room-temperature gas cell. The weak band system (ϵ ≈ 20 M^{−1} cm^{−1}) in this region is due to the S_{1}(n, π*) ← S_{0} electronic transition. The origin band was observed at 26,081(1) cm^{−1} for the undeuterated molecule and at 26,076(1) cm^{−1} for 2CHO-2,6,6-d_{3}. For the d_{0} isotopomer, about 40 vibronic transitions have been assigned in a region from −300 to +700 cm^{−1} relative to the origin band. Nearly every corresponding assignment was made for the d_{3} species. Several fundamental vibrational frequencies in the S_{1} state, as well as the five lowest ring-puckering (or inversion) energy levels in the S_{1} state, have been determined for the d_{0}/d_{3} isotopomers. The spectroscopic results are summarized below (frequencies in cm^{−1}, uncertainties ±0.5 cm^{−1}), along with results of a DFT calculation of the d_{0} isotopomer:

Vibrational frequencies of 2CHO in its S_{1} state

mode  description  d_{0} d_{0} (DFT calc) d_{3}  v′_{39} d_{0}  d_{3}
ν′_{39}  inversion  122.1  120.8  114.4 1 122.1 114.4
ν′_{38}  ring bending  251.9  249.9  236.9 2 243.8 228.6
ν′_{37}  C=C twisting  303.3  298.4  294.6 3 364.5 341.8
ν′_{36} carbonyl deformation 343.9  341.9  332.0 4 485.3 455.3
5 603.6 565.7

The inversion-level spacings in the S_{1} state indicate a barrier to planarity that is significantly higher than the 2000-cm^{−1} barrier height of the ground electronic state. Work is in progress to fit an S_{1} inversion potential to the spectroscopic data.